# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ Abstract
Large Model (LM) agents, powered by large foundation models such as GPT-4 and DALL-E 2, represent a significant step towards achieving Artificial General Intelligence (AGI). LM agents exhibit key characteristics of autonomy, embodiment, and connectivity, allowing them to operate across physical, virtual, and mixed-reality environments while interacting seamlessly with humans, other agents, and their surroundings. This paper provides a comprehensive survey of the state-of-the-art in LM agents, focusing on the architecture, cooperation paradigms, security, privacy, and future prospects. Specifically, we first explore the foundational principles of LM agents, including general architecture, key components, enabling technologies, and modern applications. Then, we discuss practical collaboration paradigms from data, computation, and knowledge perspectives towards connected intelligence of LM agents. Furthermore, we systematically analyze the security vulnerabilities and privacy breaches associated with LM agents, particularly in multi-agent settings. We also explore their underlying mechanisms and review existing and potential countermeasures. Finally, we outline future research directions for building robust and secure LM agent ecosystems.
@@
Large model, AI agents, embodied intelligence, multi-agent collaboration, security, privacy.
@@ I Introduction
I-A Background of Large Model Agents
@@
In the 1950s, Alan Turing introduced the famous Turing Test to assess whether machines could exhibit intelligence comparable to that of humans, which laid the foundation in the evolution of Artificial Intelligence (AI). These artificial entities, commonly known as “agents”, serve as the core components of AI systems. Generally, AI agents are autonomous entities capable of understanding and responding to human inputs, perceiving their environment, making decisions, and taking actions in physical, virtual, or mixed-reality settings to achieve their goals [1]. AI agents range from simple bots that follow predefined rules to complex and autonomous entities that learn and adapt through experience [2]. They can be software-based or physical entities, functioning independently or in collaboration with humans or other agents.
Since the mid-20th century, significant progress has been made in the development of AI agents [3, 4, 5], such as Deep Blue, AlphaGo, and AlphaZero, as shown in Fig. 1. Despite these advances, prior research primarily concentrated on refining specialized abilities such as symbolic reasoning or excelling in certain tasks such as Go or Chess, often neglecting the cultivation of general-purpose capabilities within AI models such as long-term planning, multi-task generalization, and knowledge retention. The challenge of creating AI agents that can flexibly adapt to a broad range of tasks and complex environments remains largely unexplored. To push the boundaries of AI agents further, it is crucial to develop powerful foundational models that integrate these critical attributes, offering a versatile basis for next-generation AI agents.
Figure 1: Evolution history of AI Agents. 1) Initial stage: Early AI research primarily focused on logical reasoning and rule-based AI agents. 2) Machine Learning (ML) agent stage: ML including supervised and unsupervised learning advanced the progress of AI agents. In 1997, IBM’s Deep Blue defeated the world chess champion. 3) Deep Learning (DL) agent stage: The combination of DL and big data significantly improved AI performance. In 2016, Deepmind’s AlphaGo defeated Go world champion Lee Sedol. 4) Large Model (LM) agent era: Transformer-based LMs such as OpenAI’s ChatGPT and GPT-4 revolutionized AI agents, ushering the era of LM agents and bringing us closer to AGI.
With the rise of Large Models (LMs), also known as large foundation models, such as OpenAI GPT-4o, Google PaLM 2, and Microsoft Copilot, LMs open up new possibilities in comprehensively enhancing the inherent capabilities of AI agents [6, 7]. As illustrated in Fig. 2, an LM agent, either in software or embodied form, generally consist of four key components: planning, action, memory, and interaction. These agents can seamlessly operate within physical, virtual, or mixed-reality environments [8, 1, 9, 10]. Particularly, LMs serve as the “brain” of AI agents and empower them with powerful capabilities in human-machine interaction (HMI), complex pattern recognition, knowledge retention, reasoning, long-term planning, generalization, and adaptability [9]. Moreover, via advanced reasoning and few/zero-shot planning techniques such as Chain-of-Thought (CoT) [11], Tree-of-Thought (ToT) [12], and reflection [13], LM agents can form intricate logical connections, enabling them to solve complex, multifaceted tasks effectively. For example, AutoGPT [14], a promising LLM agent prototype, can decompose complex tasks into several manageable sub-tasks, facilitating structured and efficient problem-solving. Integrating LMs with Retrieval-Augmented Generation (RAG) technologies [15] further allows agents to access external knowledge sources and enhance the accuracy of their responses based on retrieved information. Besides, LM agents can flexibly integrate a range of LMs, including Large Language Model (LLM) and Large Vision Model (LVM), to enable multifaceted capabilities.
@@
LM agents are recognized as a significant step towards achieving Artificial General Intelligence (AGI) and have been widely applied across fields such as web search [16], recommendation systems [17], virtual assistants [18, 19], Metaverse gaming [20], robotics [21], autonomous vehicles [22], and Electronic Design Automation (EDA) [23]. As reported by MarketsandMarkets [24], the worldwide market for autonomous AI and autonomous agents was valued at USD 48 billion in 2023 and is projected to grow at a CAGR of 43%, reaching USD 28.5 billion by 2028. LM agents have attracted global attention, and leading technology giants including Google, OpenAI, Microsoft, IBM, AWS, Oracle, NVIDIA, and Baidu are venturing into the LM agent industry.
@@
Figure 2: Use cases of LM agents. (a) a software-form LM agent acting as the virtual assistant. (b) an embodied LM agent serving as the home cleaner.
@@
Figure 3: Overview of LM agents. Each LM agent includes two parts: (i) the digital brain located in the cyberspace powered by LMs such as GPT-4o, PaLM 2, and Copilot; and (ii) the physical body such as autonomous vehicle, robot dog, and drone. Within each LM agent, the digital brain synchronizes with its physical body via intra-agent communications. LM agents communicate with one another in the cloud to share information and knowledge via inter-agent communications, establishing a network of interconnected intelligence. Each LM agent can dynamically interact with other agents, virtual/real environments, and humans. The brain of each LM agent can be deployed either as a standalone entity or in a hierarchical manner across various platforms such as cloud servers, edge devices, or end devices.
I-B Roadmap and Key Characteristics of LM Agents
Fig. 3 illustrates a future vision of LM agents, characterized by three key attributes: autonomous, embodied, and connected, paving the way toward AGI.
1) Autonomous Intelligence. Autonomous intelligence in LM agents refers to their ability to operate independently, making proactive decisions without continuous human input. As depicted in Fig. 2(a), an LM agent can maintain an internal memory that accumulates knowledge over time to guide future decisions and actions, enabling continuous learning and adaptation in dynamic environments [25]. Additionally, LM agents can autonomously utilize a variety of tools (e.g., search engines and external APIs) to gather information or or create new tools to handle intricate tasks [26]. By collaborating or competing with humans or other agents, LM agents can effectively enhance their decision-making capabilities [27].
2) Embodied Intelligence. Despite recent advancements, LMs typically passively respond to human commands in the text, image, or multimodal domain, without engaging directly with the physical world [7]. Embodied agents, on the other hand, can actively perceive and act upon their environment, whether digital, robotic, or physical, using sensors and actuators [25, 21]. The shift to LM-empowered agents involves creating embodied AI systems capable of understanding, learning, and solving real-world challenges. As depicted in Fig. 2(b), LM agents actively interact with environments and adapt actions based on real-time feedbacks. For example, a household robot LM agent tasked with cleaning can generate tailored strategies by analyzing the room layout, surface types, and obstacles, instead of merely following generic instructions.
3) Connected Intelligence. Connected LM agents extend beyond the capabilities of individual agents, playing a critical role in tackling complex, real-world tasks [28]. For example, in autonomous driving, connected autonomous vehicles, serving as LM agents, share real-time sensory data, coordinate movements, and negotiate passage at intersections to optimize traffic flow and enhance safety. As depicted in Fig. 3, by interconnecting numerous LM agents into the Internet of LM agents, connected LM agents can freely share sensory data and task-oriented knowledge. By fully harnessing the computational power of various specialized LMs, it fosters cooperative decision-making and collective intelligence. Theereby, the collaboration across data, computation, and knowledge domains enhances individual agent performance and adaptability. Additionally, these interactions enable LM agents to form social connections and attributes, contributing to the development of an agent society [29, 30].
I-C Motivation for Securing Connected LM Agents
@@
Despite the bright future of LM agents, security and privacy concerns remain significant barriers to their widespread adoption. Throughout the life-cycle of LM agents, numerous vulnerabilities can emerge, ranging from adversarial examples [31], agent poisoning [32], LM hallucination [33], to pervasive data collection and memorization [34].
@@
1) Security vulnerabilities. LM agents are prone to “hallucinations”, where their foundational LMs generate plausible but incorrect outputs not grounded in reality [33]. In multi-agent settings, the hallucination phenomenon can propagate misinformation, compromise decision-making, cause task failures, and pose risks to both physical entities and human. Moreover, maintaining the integrity and authenticity of sensory data and prompts used in training and inference is crucial, as biased or compromised inputs can lead to inaccurate or unfair outcomes [35]. Attacks such as adversarial manipulations [31], poisoning [36], and backdoors [37] further threaten LM agents by allowing malicious actors to manipulate inputs and deceive the models. In collaborative environments, agent poisoning behaviors [32], where malicious agents disrupt the behavior of others, can undermine the collaborative systems. Additionally, integrating LM agents into Cyber-Physical-Social Systems (CPSS) expands the attack surface, enabling adversaries to exploit vulnerabilities within interconnected systems.
@@
2) Privacy breaches. LM agents’ extensive data collection and memorization processes raise severe risks of data breaches and unauthorized access. These agents often handle vast amounts of personal and sensitive business information for both To-Customer (ToC) and To-Business (ToB) applications, raising concerns about data storage, processing, sharing, and control [38]. Additionally, LMs can inadvertently memorize sensitive details from their training data, potentially exposing private information during interactions [34]. Privacy risks are further compounded in multi-agent collaborations, where LM agents might inadvertently leak sensitive information about users, other agents, or their internal operations during communication and task execution.
I-D Related Surveys and Contributions
@@
Recently, LM agents have garnered significant interest across academia and industry, leading to a variety of research exploring their potential from multiple perspectives. Notable survey papers in this field are as below. Andreas et al. [29] present a toy experiment for AI agent construction and case studies on modeling communicative intentions, beliefs, and desires. Wang et al. [39] identify key components of LLM-based autonomous agents (i.e., profile, memory, planning, and action) and the subjective and objective evaluation metrics. Besides, they discuss the applications of LLM agents in engineering, natural science, and social science. Xi et al. [9] present a general framework for LLM agents consisting of brain, action, and perception. Besides, they explore applications in single-agent, multi-agent, and human-agent collaborations, as well as agent societies. Zhao et al. [2] offer a systematic review of LLMs in terms of pre-training, adaptation tuning, utilization, and capacity assessment. Besides, background information, mainstream technologies, and critical applications of LLMs are introduced. Xu et al. [40] provide a tutorial on key concepts, architecture, and metrics of edge-cloud AI-Generated Content (AIGC) services in mobile networks, and identify several use cases and implementation challenges. Huang et al. [1] offer a taxonomy of AI agents in virtual/physical environments, discuss cognitive aspects of AI Agents, and survey the applications of AI agents in robotics, healthcare, and gaming. Cheng et al. [10] review key components of LLM agents (including planning, memory, action, environment, and rethinking) and their potential applications. Planning types, multi-role relationships, and communication methods in multi-agent systems are also reviewed. Masterman et al. [8] provide an overview of single-agent and multi-agent architectures in industrial projects and present the insights and limitations of existing research. Guo et al. [41] discuss the four components (i.e., interface, profiling, communication, and capabilities acquisition) of LLM-based multi-agent systems and present two lines of applications in terms of problem solving and world simulation. Durante et al. [42] introduce multimodal LM agents and a training framework including learning, action, congnition, memory, action, and perception. They also discuss the different roles of agents (e.g., embodied, simulation, and knowledge inference), as well as the potentials and experimental results in different applications including gaming, robotics, healthcare, multimodal tasks, and Natural Language Processing (NLP). Hu et al. [20] outline six key components (i.e., perception, thinking, memory, learning, action, and role-playing) of LLM-based game agents and review existing LLM-based game agents in six types of games. Xu et al. [43] provide a comprehensive survey of enabling architectures and challenges for LM agents in gaming. Qu et al. [44] provide a comprehensive survey on integrating mobile edge intelligence (MEI) with LLMs, emphasizing key applications of deploying LLMs at the network edge along with state-of-the-art techniques in edge LLM caching, delivery, training, and inference.
@@
Existing survey works on LM agents mainly focus on the general framework design for single LLM agents and multi-agent systems and their potentials in specific applications. Distinguished from the above-mentioned existing surveys, this survey focuses on the networking aspect of LM agents, including the general architecture, enabling technologies, and collaboration paradigms to construct networked systems of LM agents in physical, virtual, or mixed-reality environments. Moreover, with the advances of LM agents, it is urgent to study their security and privacy in future AI agent systems. This work comprehensively reviews the security and privacy issues of LM agents and discusses the existing and potential defense mechanisms, which are overlooked in existing surveys. Table I compares the contributions of our survey with previous related surveys in the field of LM agents.
@@
Discuss key components of LLM-based multi-agent system
@@
architecture, enabling technologies, networking modes, and
collaboration paradigms of connected LM agents, discussions
on security/privacy threats, state-of-the-art solutions, and
open research issues in connected LM agent design.
@@
In this paper, we present a systematic review of the state-of-the-arts in both single and connected LM agents, focusing on security and privacy threats, existing and potential countermeasures, and future trends. Our survey aims to 1) provide a broader understanding of how LM agents work and how they interact in multi-agent scenarios, 2) examine the scope and impact of security and privacy challenges associated with LM agents and their interactions, and 3) highlight effective strategies and solutions for defending against these threats to safeguard LM agents in various intelligent applications. The main contributions of this work are four-fold.
  • We comprehensively review recent advances in LM agent construction across academia and industry. We investigate the working principles of LM agents, including the general architecture, key components (i.e., planning, memory, action, interaction, and security modules), and enabling technologies. The industrial prototypes and potential applications of LM agents are also discussed.
  • We systematically categorize interaction patterns for LM agents (i.e., agent-agent, agent-human, and agent-environment interactions) and their interaction types (i.e., cooperation, partial cooperation, and competition). We explore practical collaboration paradigms of LM agents from the aspects of data cooperation, computation cooperation, and knowledge cooperation.
  • We comprehensively analyze existing and potential security and privacy threats, their underlying mechanisms, categorization, and challenges for both single and connected LM agents. We also review state-of-the-art countermeasures and examine their feasibility in securing LM agents.
  • Lastly, we discuss open research issues and point out future research directions from the perspectives of energy-efficient and green LM agents, fair and explainable LM agents, cyber-physical-social secure agent systems, value networks of agent ecosystem, aiming to advance the efficiency and security of LM agents.

### v2

# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

@@ Abstract
With the rapid advancement of large models (LMs), the development of general-purpose intelligent agents powered by LMs has become a reality. It is foreseeable that in the near future, LM-driven general AI agents will serve as essential tools in production tasks, capable of autonomous communication and collaboration without human intervention. This paper investigates scenarios involving the autonomous collaboration of future LM agents. We review the current state of LM agents, the key technologies enabling LM agent collaboration, and the security and privacy challenges they face during cooperative operations. To this end, we first explore the foundational principles of LM agents, including their general architecture, key components, enabling technologies, and modern applications. We then discuss practical collaboration paradigms from data, computation, and knowledge perspectives to achieve connected intelligence among LM agents. After that, we analyze the security vulnerabilities and privacy risks associated with LM agents, particularly in multi-agent settings, examining underlying mechanisms and reviewing current and potential countermeasures. Lastly, we propose future research directions for building robust and secure LM agent ecosystems.
@@
Large models, AI agents, embodied intelligence, networking, multi-agent collaboration, security, privacy.
@@ I Introduction
I-A Background of Large Model Based Agents
@@
In the 1950s, Alan Turing introduced the Turing Test to assess whether machines could exhibit intelligence comparable to humans. These artificial entities, commonly known as “agents”, serve as the core components of artificial intelligence (AI) systems. AI agents (also known as agentic AI [1]) are autonomous entities capable of understanding and responding to human inputs, perceiving their environment, making decisions, and taking actions across physical, virtual, or mixed-reality settings to achieve specific goals [2]. They can be software-based or physical entities, functioning independently or in collaboration with humans or other agents. Since the mid-20th century, significant progress has been made in the development of AI agents [3, 4], such as Deep Blue, AlphaGo, and AlphaZero, as shown in Fig. 1. Despite these advances, prior research primarily concentrated on refining specialized abilities such as symbolic reasoning or excelling in certain tasks such as Go or Chess, often neglecting the cultivation of general-purpose capabilities within AI models such as long-term planning, multi-task generalization, and knowledge retention.
Figure 1: Evolution history of AI agents.
With the rise of large models (LMs) [5], including large language models (LLMs) such as OpenAI GPT-4, Google PaLM 2, and Microsoft Copilot, as well as large vision models (LVMs) such as DALL-E 2, LM-based agents (also known as LM agents or agentic LMs [6]) are unlocking new possibilities [7]. LM agents significantly enhance the inherent capabilities of AI systems, providing a versatile foundation for the next-generation AI agents [8]. Serving as the “brain” of AI agents, LMs empower them with advanced capabilities in human-machine interaction (HMI), few/zero-shot planning, contextual understanding, knowledge retention, and general-purpose task solving across physical, virtual, or mixed-reality environments [9, 10, 11]. LM agents generally fall into two categories:
  • Software-form LM agents, such as AutoGPT [12] and AutoGen [13], can autonomously interpret human instructions and use various tools (e.g., search engines and external APIs) to gather information and complete intricate tasks [14]. For instance, as shown in Fig. 2(a), an LM-powered virtual assistant can generate personalized travel plans, set reminders, and manage tasks while continuously learning and adapting in dynamic environments.
  • Embodied LM agents, such as FigureAI’s Figure 02 and Tesla’s Optimus, engage directly with the physical world. These agents perceive and interact with their surroundings, allowing them to solve real-world problems [15]. For instance, as shown in Fig. 2(b), an LM-powered household robot analyzes room layouts, surface types, and obstacles, to devise customized cleaning strategies rather than merely following generic instructions.
@@
LM agents are recognized as a significant step towards achieving artificial general intelligence (AGI) and have been widely applied across fields such as web search [16], recommendation systems [17], virtual assistants [18, 19], Metaverse gaming [20], robotics [15], autonomous vehicles [21], and automated attack penetration [22]. As reported by MarketsandMarkets [23], the worldwide market for autonomous AI and autonomous agents was valued at USD 4.8 billion in 2023 and is projected to grow at a CAGR of 43%, reaching USD 28.5 billion by 2028. LM agents have attracted global attention, and leading technology giants including Google, OpenAI, Microsoft, IBM, AWS, Oracle, NVIDIA, and Baidu are venturing into the LM agent industry.
@@
Figure 2: Use cases of LM agents: (a) a software-form LM agent acting as the virtual assistant; (b) an embodied LM agent serving as the home cleaner. An LM agent, either in software or embodied form, generally consists of four key components: interaction, planning, action, and memory.
@@
Figure 3: Overview of LM agent networks. Each LM agent includes two parts: (i) the brain located in the cyberspace powered by LMs such as GPT-4o, PaLM 2, and Copilot; and (ii) the physical or digital body such as autonomous vehicle, robot dog, UAV, and digital human. Within each LM agent, the cyber brain synchronizes with its physical/digital body via intra-agent communications. LM agents communicate with one another in the cloud to share task-oriented, knowledge, and inference results via inter-agent communications, establishing a network of interconnected intelligence. Each LM agent dynamically interact with other agents, virtual/real environments, and humans. The brain of each LM agent can be deployed either as a standalone entity or in a hierarchical manner across various platforms such as cloud servers, edge servers, or end devices.
I-B Motivation of LM Agent Networks
Beyond the intelligence of single LM agents, connected LM agents can form LM agent networks to cooperatively address complex tasks beyond the capacity of individual agents [24]. As depicted in Fig. 3, within LM agent networks, connected LM agents can freely share sensory data, task-oriented knowledge, and LM inference results. By leveraging collective computational power and shared expertise from specialized LM agents, it fosters cooperative decision-making and collective intelligence. For instance, in autonomous driving, connected autonomous vehicles, acting as LM agents, share real-time sensory data, coordinate movements, and negotiate passage at intersections to optimize traffic flow and enhance safety. Each LM agent consists of an LM-powered brain (hosted in the cloud, edge, or end devices) and a physical or digital body. For embodied LM agents, a tiny local LM within the physical body handles local inference, while compute-intensive tasks are offloaded to the cloud LM, enabling collaborative computation between the cloud brain and the local brain. Each LM agent can interact with other agents, virtual/real environments, and humans. Communication within LM agent networks occurs via two primary modes: (i) intra-agent communications for seamless synchronization of status, data, and knowledge between the brain and body to enable real-time functionality; and (ii) inter-agent communications for efficient information exchange and computational coordination among multiple LM agents.
LM agent networks represent a transformative leap in intelligent systems, which incorporate a range of cutting-edge technologies as its foundation. Particularly, foundation models enable cognitive capabilities through advanced reasoning and planning, knowledge-related technologies integrate external knowledge sources for context-aware actions, HMI technologies ensure seamless human-agent interactions via natural language processing (NLP) and augmented/virtual/mixed reality (AR/VR/MR), digital twin technologies synchronize physical and digital states through intra-agent communications, and multi-agent collaboration technologies foster efficient teamwork via inter-agent communications. However, their full potential faces key challenges, including but not limited:
  • Dynamic Construction of LM Agent Networks. Constructing dynamic LM agent networks requires addressing the versatility-efficiency-portability trilemma [25]. The network should support diverse tasks and applications (versatility), achieve high resource utilization with low cost (efficiency), and operate across heterogeneous platforms and environments (portability). The inherent spatiotemporal dynamics introduce additional complexity. From temporal perspectives, agents’ roles and behaviors evolve over time, necessitating adaptive reconfiguration mechanisms to ensure continuity in long-term operations. From spatial perspectives, the topologies of LM agent networks change due to mobility and environmental changes, requiring robust coordination protocols to maintain network coherence.
  • Collaborative LM Service Provisioning in Heterogeneous Networks. Once constructed, LM agent networks should enable heterogeneous agents to collaboratively provide LM services. Given the substantial computational resources required to run a complete LM, edge or terminal agents often lack the capacity to operate a full-scale LM. There exist two primary strategies: (i) LM lightweighting, e.g., knowledge distillation [26], quantization [27], pruning [28], and hardware acceleration [29], for deploying compact LM variants tailored to individual agents’ computational capacities. Collaboration is facilitated through methods such as role-playing for task-specific role assignment and distributed consensus for coordinated decision-making. (ii) LM sharding, e.g., split learning (SL) [30] and mixture-of-experts (MoE) [31], for distributed LM reconstruction, enabling agents to share workloads efficiently. The full-scale LM is segmented into smaller, manageable shards distributed across agents, allowing each edge or terminal agent to handle a portion of the model that aligns with its resource capabilities.
  • Autonomous Optimization and Secure Collaboration. Dynamic resources (e.g., data, knowledge, and computational power) allocation is vital to maximize throughput, minimize latency, and maintain seamless communication across cloud-edge-end layers. Security mechanisms in collaboration are also crucial to ensure trustworthiness (i.e., LM agents operate securely and transparently in collaborative tasks) and privacy preservation (i.e., safeguarding sensitive data during agent interactions).
I-C Motivation of Securing LM Agent Networks
@@
Despite the promising future of connected LM agents, security and privacy concerns pose significant barriers to their widespread adoption. Throughout the life-cycle of LM agents, numerous vulnerabilities can emerge, ranging from adversarial examples [32], agent poisoning [33], LM hallucination [34], to pervasive data collection and memorization risks [35].
@@
1) Security vulnerabilities. LM agents are susceptible to “hallucinations”, where their LMs generate outputs that are plausible but incorrect and not grounded in reality, potentially spreading misinformation in multi-agent environments [34]. Hallucinations can compromise the reliability of decision-making, cause task failures, and pose risks to both physical entities and humans. Ensuring the integrity of intermediate inference results in collaborative systems is crucial, as biased or manipulated inputs can lead to unfair or erroneous outcomes [36]. Attacks such as adversarial manipulations [32], poisoning [37], and backdoors [38] exacerbate these vulnerabilities by allowing malicious actors to manipulate inputs and deceive the models. In collaborative environments, agent poisoning behaviors [33], where malicious agents disrupt the operation of others, can undermine the collaborative systems. Additionally, integrating LM agents into cyber-physical-social systems (CPSS) expands the attack surface, providing more opportunities for malicious exploitation within interconnected systems.
@@
2) Privacy breaches. The extensive data collection processes and data memorization capabilities of LM agents raise significant risks related to data breaches and unauthorized access. These agents often process large volumes of personal and sensitive business information in both to-customer (ToC) and to-business (ToB) applications, which raises concerns regarding data storage, processing, and sharing [39]. Furthermore, LMs can inadvertently memorize sensitive details from training data and past interactions, potentially exposing private information during subsequent interactions [35]. These privacy risks are exacerbated in multi-agent collaborations, where LM agents may unintentionally leak sensitive information about users, other agents, or their internal operations during communication and task execution.
I-D Related Surveys and Our Contributions
@@
Recently, LM agents have garnered significant interest across academia and industry, leading to a variety of research exploring their potential from multiple perspectives. Notable survey papers in this field are as below. Andreas et al. [40] present a toy experiment for AI agent construction and case studies on modeling communicative intentions, beliefs, and desires. Wang et al. [41] identify key components of LLM-based autonomous agents (i.e., profile, memory, planning, and action) and the subjective and objective evaluation metrics. Besides, they discuss the applications of LLM agents in engineering, natural science, and social science. Xi et al. [10] present a general framework for LLM agents consisting of brain, action, and perception. Besides, they explore applications in single-agent, multi-agent, and human-agent collaborations, as well as agent societies. Zhao et al. [42] offer a systematic review of LLMs in terms of pre-training, adaptation tuning, utilization, and capacity assessment. Besides, background information, mainstream technologies, and critical applications of LLMs are introduced. Xu et al. [43] provide a tutorial on key concepts, architecture, and metrics of edge-cloud AI-generated content (AIGC) services in mobile networks, and identify several use cases and implementation challenges. Huang et al. [2] offer a taxonomy of AI agents in virtual/physical environments, discuss cognitive aspects of AI Agents, and survey the applications of AI agents in robotics, healthcare, and gaming. Cheng et al. [11] review key components of LLM agents (including planning, memory, action, environment, and rethinking) and their potential applications. Planning types, multi-role relationships, and communication methods in multi-agent systems are also reviewed. Masterman et al. [9] provide an overview of single-agent and multi-agent architectures in industrial projects and present the insights and limitations of existing research. Guo et al. [44] discuss the four components (i.e., interface, profiling, communication, and capabilities acquisition) of LLM-based multi-agent systems and present two lines of applications in terms of problem solving and world simulation. Durante et al. [45] introduce multimodal LM agents and a training framework including learning, action, congnition, memory, action, and perception. They also discuss the different roles of agents (e.g., embodied, simulation, and knowledge inference), as well as the potentials and experimental results in different applications including gaming, robotics, healthcare, multimodal tasks, and NLP. Hu et al. [20] outline six key components (i.e., perception, thinking, memory, learning, action, and role-playing) of LLM-based game agents and review existing LLM-based game agents in six types of games. Xu et al. [5] provide a comprehensive survey of enabling architectures and challenges for LM agents in gaming. Qu et al. [46] provide a comprehensive survey on integrating mobile edge intelligence (MEI) with LLMs, emphasizing key applications of deploying LLMs at the network edge along with state-of-the-art techniques in edge LLM caching, delivery, training, and inference.
@@
Existing surveys on LM agents primarily focus on the general framework design of LLM agents and multi-agent systems, as well as their potential in specific applications. Distinguished from the above-mentioned existing surveys, this survey focuses on the networking aspect of LM agents, including the general architecture, enabling technologies, key characteristics, and collaboration paradigms to construct networked systems of LM agents within physical, virtual, or mixed-reality environments. Additionally, with the advances of LM agents, it is urgent to examine their security and privacy in future AI agent systems. This work comprehensively reviews the security and privacy issues of LM agents and discusses both existing and potential defense mechanisms, which are overlooked in previous surveys. Table I compares the contributions of our survey with previous related surveys in the field of LM agents.
@@
Discuss key components of LLM-based multi-agent systems
@@
architecture, enabling technologies, key characteristics, and
cooperation paradigms of connected LM agents, discussions
on security/privacy threats, state-of-the-art countermeasures,
and open research issues in future LM agent systems.
@@
In this paper, we present a systematic review of the state-of-the-arts in both single and connected LM agents, focusing on cooperation paradigms, security and privacy threats, existing and potential countermeasures, and future trends. Our survey aims to 1) provide a broader understanding of how LM agents work and how they cooperate in multi-agent scenarios, 2) examine the scope and impact of security and privacy challenges associated with LM agents and their interactions, and 3) highlight effective strategies and solutions for defending against these threats to safeguard LM agents in various intelligent applications. The main contributions of this work are four-fold.
  • We comprehensively review recent advances in LM agent construction across academia and industry. We investigate the fundamentals of LM agents, including the general architecture, and key components (i.e., planning, memory, action, interaction, and security modules). Industrial prototypes and potential applications of LM agents are also discussed.
  • We discuss key characteristics and enabling technologies of LM agent networks and explore practical cloud-edge-end collaboration paradigms from the aspects of data cooperation, computation cooperation, and knowledge cooperation. We systematically discuss interaction strategies of LM agents (i.e., cooperation, partial cooperation, and competition).
  • We comprehensively analyze existing and potential security and privacy threats, their underlying mechanisms, categorization, and challenges for both single and connected LM agents. We also review state-of-the-art countermeasures and examine their feasibility in securing LM agent networks.
  • Lastly, we discuss open research opportunities and point out future research directions of LM agent networks, aiming to inspire more ongoing research and innovations in this field.